# for Large Language Models

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have advanced graph structure understanding via recursive information exchange and aggregation among graph nodes. To improve model robustness, self-supervised learning (SSL) has emerged as a promising approach for data augmentation. However, existing methods for generating pre-trained graph embeddings often rely on fine-tuning with specific downstream task labels, which limits their usability in scenarios where labeled data is scarce or unavailable. To address this, our research focuses on advancing the generalization capabilities of graph models in challenging zero-shot learning scenarios. Inspired by the success of large language models (LLMs), we aim to develop a graph-oriented LLM that can achieve high generalization across diverse downstream datasets and tasks, even without any information available from the downstream graph data. In this work, we present the GraphGPT framework that aligns LLMs with graph structural knowledge with a graph instruction tuning paradigm. Our framework incorporates a text-graph grounding component to establish a connection between textual information and graph structures. Additionally, we propose a dual-stage instruction tuning paradigm, accompanied by a lightweight graph-text alignment projector. This paradigm explores self-supervised graph structural signals and task-specific graph instructions, to guide LLMs in understanding complex graph structures and improving their adaptability across different downstream tasks. Our framework is evaluated on supervised and zero-shot graph learning tasks, demonstrating superior generalization and outperforming state-of-the-art baselines.
*** Corresponding author, chaohuang75@gmail.com
copyright: none
@@ 2. Preliminaries
Graph-structured Data. represents information as a collection of entities (nodes) and the relationships (edges) between them. A graph is formally denoted as $\mathcal{G}(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$, encompassing several core components. The node set $\mathcal{V}$ represents a collection of nodes, with $|\mathcal{V}|=N$ indicating the total number of nodes in the graph. The edge set $\mathcal{E}$ characterizes the relationships or connections between the nodes. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ encodes the topology of the graph, with each element $A_{i,j}$ indicating the presence or absence of an edge between nodes $i$ and $j$. The feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$ contains the attribute or feature information associated with each node, where $F$ represents the dimensionality of features.
Graph Neural Networks. have emerged as a powerful framework for representation learning from graph-structured data. Unlike traditional neural networks that operate on grid-like data, GNNs can effectively capture and model the complex relationships and dependencies present in graphs. Specifically, GNNs leverage the inherent structure of the graph, consisting of nodes and edges, to learn expressive node representations through iterative message propagation and aggregation operations, presented as follows.
@@
After applying the message passing and aggregation mechanism in Graph Neural Networks (GNNs), the encoded feature vector of node $v$ at the $l$-th layer is denoted as $h_{v}^{(l)}$. The $\text{Propagate}^{(l)}$ function performs message passing by aggregating information from the neighboring nodes of $v$ at the $l$-th layer. The $\text{Aggregate}^{(l)}$ function then combines the aggregated information with the previous layer’s representation of node $v$ to generate the updated representation $h_{v}^{(l)}$. By encoding graph structural information with the learned representations, GNNs can be customized for various downstream graph learning tasks, such as node classification and link prediction.
@@ 3.1. Structural Information Encoding with Text-Graph Grounding
To enhance the understanding of graph structural information by large language models, our framework emphasizes aligning the encoding of graph structures with the natural language space. This alignment aims to enable language models to effectively comprehend and interpret the structural elements of the graph, leveraging their inherent language understanding capabilities. To achieve this objective, we introduce a text-graph grounding paradigm that generates prompts designed to preserve the graph’s structural context for language models. This paradigm acts as a bridge, connecting the semantic understanding of textual information with the inherent structural relationships found within the graph.
@@
Text-Structure Alignment. To align graph structure information within Language Models (LLMs) more effectively, our main focus is to explore the encoding of graph structures that can collaborate well with LLMs. Inspired by prior works (Radford et al., 2021; Wen and Fang, 2023), we incorporate textual information into the graph structure encoding process in a contrastive manner. In our approach, we directly integrate a graph encoder with pre-trained parameters into our GraphGPT framework, enabling seamless integration of the graph encoder’s capabilities in our framework. Formally, let $\mathcal{G}(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$ represent a graph with raw textual contents $\mathbf{C}=\{c_{i}\in\mathbb{R}^{l_{i}\times d},1\leq i\leq N\}$ for $N$ nodes, where $l_{i}$ denotes the length of the textual content for the $i$-th node. We obtain encoded graph representations $\hat{\mathbf{H}}\in\mathbb{R}^{N\times d}$ and encoded text representations $\hat{\mathbf{T}}\in\mathbb{R}^{N\times d}$ as follows:
@@
We utilize the graph encoder, denoted as $f_{\mathbf{G}}$, to take the graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$ as input and generates structure-level graph representations. We employ a text encoder, such as a transformer or Bert, denoted as $f_{\mathbf{T}}$, to encode the raw textual contents $\mathbf{C}$ associated with the nodes. This step produces encoded text representations of nodes. We further apply row-wise L2 normalization using the norm function. Formally, the text-structure alignment with the cross-modalities is conducted as follows:
$\Gamma_{i}$ $=(g_{i}^{(1)}(\hat{\mathbf{H}})g_{i}^{(2)}(\hat{\mathbf{T}})^{\top})\cdot\exp(\tau)$
(4) $\mathcal{L}$ $=\sum_{i}\frac{1}{2}\lambda_{i}(\text{CE}(\Gamma_{i},\mathbf{y})+\text{CE}(\Gamma_{i}^{\top},\mathbf{y}))$
In our text-graph grounding, we use the label $\mathbf{y}=(0,1,\cdots,n-1)^{\top}$ for the contrastive alignment objective. The transformation functions for different grounding designs are denoted as $g_{i}^{(1)}$ and $g_{i}^{(2)}$. We employ a graph transformer (Yun et al., 2019b) as the graph encoder and a vanilla transformer (Vaswani et al., 2017) as the text encoder.
Figure 3. Workflow of text-structure alignment.
@@ 3.2.1. Self-Supervised Instruction Tuning
In the first stage of our graph instruction tuning paradigm, we introduce self-supervised instruction tuning. This mechanism allows us to inject graph domain-specific structural knowledge into the language model, improving its reasoning abilities and enabling it to effectively comprehend the contextual information embedded in the graph’s structure. To achieve this, we incorporate self-supervised signals derived from unlabeled graph structures as instructions for model tuning. In particular, we design a structure-aware graph matching task that guides the language model in distinguishing between different graph tokens using natural language tokens. This instruction task plays a crucial role in accurately associating graph tokens with their corresponding textual descriptions, thereby deepening the model’s understanding of the graph with the provided guidance.
Instruction Design. The instruction for our graph matching task consists of three components: i) graph information, ii) human question, and iii) GraphGPT response. In this task, we treat each node in the graph as a central node and perform h-hops with random neighbor sampling, resulting in a subgraph structure. The natural language input for the LLM is the human question. In the context of the graph matching task, the instruction includes the indicator token <graph> and a shuffled list of node text information. For example, in a citation graph, the node text information corresponds to paper titles. The objective of the LLM in the graph matching task is to align each graph token with its corresponding node text information. This requires reordering the node text information list based on the sequence of graph tokens, effectively associating each graph token with its relevant textual description.
@@ 3.2.2. Task-Specific Instruction Tuning
In the second stage, we propose task-specific instruction tuning. This step is designed to customize the model’s reasoning behavior to meet the specific constraints and requirements of different graph learning tasks, such as node classification or link prediction. By fine-tuning the LLM using task-specific graph instructions, we guide the model to generate responses that are better suited for the particular graph learning task at hand. This further improves the model’s adaptability and performance in handling various graph learning tasks.
Instruction Design. We adopt a similar instruction template, which consists of three parts. To generate graph information for each node, we employ the same neighbor sampling approach used in the first stage. This approach ensures that relevant graph information is captured, with each node acting as the central node. For the node classification task, the human question instruction contains both the indicator token <graph> and specific text information about the central node. This instruction prompts the language model to predict the category of the central node based on both the graph structure data and the accompanying text information. An example of the instruction data for different tasks can be seen in Figure 4, providing a visual representation of how the instruction is structured and presented to the language model.
@@ 3.3. Chain-of-Thought (CoT) Distillation
When faced with diverse graph data, language models may encounter new or unfamiliar patterns and structures. This distribution shift can pose challenges in generating accurate and coherent responses, especially when the number of node classes varies across different types of graph data. To address this challenge and boost accuracy in the presence of distribution shift, it is essential to equip our GraphGPT with step-by-step reasoning abilities. In this regard, we propose utilizing the Chain-of-Thought (COT) technique (Wei et al., 2022b), which explicitly models the flow of thoughts and reasoning steps. By incorporating COT, our language model improves the coherence and consistency of generated text. It enables the model to follow a logical progression of ideas, enhancing its ability to understand and reason about the given graph data.
However, incorporating the Chain-of-Thought (COT) technique can be challenging due to the influence of model parameter scale (Shridhar et al., 2023). To overcome this, we draw inspiration from previous research (Shridhar et al., 2023) and employ a distillation approach to extract valuable knowledge from a closed-source, powerful language model like ChatGPT (with over 200 billion parameters). This enables us to generate high-quality COT instructions and enhance our model’s COT reasoning capabilities while avoiding an increase in parameters.
COT Distillation Paradigm. Our approach involves designing tailored Chain-of-Thought (COT) prompts specifically for node-specific tasks. For the node classification task within a citation graph, we provide the abstract, title of the paper represented by the node, and a description of the classification task as part of the input. We then employ the GPT-3.5 language model (LLM) in our implementation to perform step-by-step reasoning. We prompt the LLM to arrive at the final answer by engaging in a sequential thought process. In the generated output, the LLM not only provides predictions for the node classes but also offers detailed explanations for each prediction. This ensures that the model’s reasoning and decision-making process are transparent and comprehensible. To further improve the performance, we integrate the generated COT instruction data with the previously designed instructions for the task-specific instruction tuning stage. With the integrated instructions, we proceed with the proposed instruction tuning paradigm.
@@ 4. Evaluation
We conduct experiments to validate the effectiveness of our framework in various settings and address key research questions.
@@ 4.1.1. Data Descriptions
We evaluate the performance of our GraphGPT using three datasets: OGB-arxiv, PubMed, and Cora. The OGB-arxiv dataset (Hu et al., 2020c) represents a directed graph that captures the citation network among computer science arXiv papers indexed by MAG (Wang et al., 2020). Each paper in the dataset is associated with a research category, manually labeled by the authors and arXiv moderators. These research categories are selected from a set of 40 subject areas. The PubMed dataset (He et al., 2023) consists of 19,717 scientific publications on diabetes obtained from the PubMed database. The publications are categorized into Experimental induced diabetes, Type 1 diabetes, and Type 2 diabetes. Additionally, the dataset includes a citation network with 44,338 links. The Cora dataset (Wen and Fang, 2023) comprises 25,120 research papers connected through citations. We utilize an expanded version of the Cora dataset, which is larger and has more classes (70 in total) compared to previous versions (Kipf and Welling, 2017).
@@ 4.1.2. Evaluation Protocols
To ensure compatibility and enable comparison across different datasets, we map the node features into the same vector space. We achieve this by encoding the raw text information using a pre-trained BERT model (Devlin et al., 2019). In our experiments, we divide the Cora and PubMed datasets into three parts: training, validation, and testing, following a ratio of 3:1:1. This partitioning scheme aligns with the approaches described in the works (He et al., 2023; Wen and Fang, 2023). For the OGB-arxiv dataset, we adhere to the public split setting (Hu et al., 2020c), which employs a ratio of 6:2:3 for training, validation, and testing. To evaluate the performance of our model, we utilize three commonly adopted evaluation metrics: Accuracy and Macro F1 for node classification, and AUC (Area Under the Curve) for link prediction.
@@ 4.1.3. Baseline Methods
In our performance comparison, we consider various state-of-the-art methods for comprehensive evaluation. (i) The first category includes MLP, which employs a Multilayer Perceptron for node representation. (ii) The second category comprises representative graph neural encoders, including GraphSAGE (Hamilton et al., 2017), GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), and RevGNN (Li et al., 2021). (iii) The third category focuses on the self-supervised approach DGI (Velickovic et al., 2019) for graph learning. (iv) The fourth category explores knowledge distillation-enhanced GNNs, with GKD (Yang et al., 2022b) and GLNN (Zhang et al., 2022b) as notable methods. (v). The fifth category showcases recently proposed strong graph transformer networks, with NodeFormer (Wu et al., 2023b) and DIFFormer (Wu et al., 2023a) as competitors. (vi) Lastly, we consider open-sourced LLMs, such as Baichuan-7B, vicuna-7B-v1.1, and vicuna-7B-v1.5 as baselines for understanding text-attributed graph data. For more detailed descriptions of the baselines, please refer to the Appendix.

### v3

# GraphGPT: Graph Instruction Tuning for Large Language Models

@@ Abstract.
Graph Neural Networks (GNNs) have evolved to understand graph structures through recursive exchanges and aggregations among nodes. To enhance robustness, self-supervised learning (SSL) has become a vital tool for data augmentation. Traditional methods often depend on fine-tuning with task-specific labels, limiting their effectiveness when labeled data is scarce. Our research tackles this by advancing graph model generalization in zero-shot learning environments. Inspired by the success of large language models (LLMs), we aim to create a graph-oriented LLM capable of exceptional generalization across various datasets and tasks without relying on downstream graph data. We introduce the GraphGPT framework, which integrates LLMs with graph structural knowledge through graph instruction tuning. This framework includes a text-graph grounding component to link textual and graph structures and a dual-stage instruction tuning approach with a lightweight graph-text alignment projector. These innovations allow LLMs to comprehend complex graph structures and enhance adaptability across diverse datasets and tasks. Our framework demonstrates superior generalization in both supervised and zero-shot graph learning tasks, surpassing existing benchmarks. The open-sourced model implementation of our GraphGPT is available at https://github.com/HKUDS/GraphGPT.
Large Language Models, Graph Learning, Instruction Tuning
copyright: nonejournalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657775isbn: 979-8-4007-0431-4/24/07ccs: Information systems Data miningccs: Mathematics of computing Graph algorithmsccs: Information systems Language models
@@ 2. Preliminaries
Graph-structured Data. represents information as entities (nodes) and the relationships (edges) between them. A graph is denoted as $\mathcal{G}(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$, comprising key components. The node set $\mathcal{V}$ represents the collection of nodes, with $|\mathcal{V}|=N$ indicating the total number of nodes. The edge set $\mathcal{E}$ characterizes the relationships between nodes. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ encodes the graph’s topology, with each element $A_{i,j}$ indicating the presence or absence of an edge between nodes $i$ and $j$. The feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$ contains attribute or feature information associated with each node, where $F$ represents the feature dimensionality.
Graph Neural Networks. have become a powerful framework for learning representations from graph-structured data. Unlike traditional neural networks that process grid-like data, GNNs excel in capturing the intricate relationships and dependencies within graphs. They utilize the graph’s structure-comprising nodes and edges-to derive expressive node representations through repeated message propagation and aggregation operations.
@@
In Graph Neural Networks, the feature vector of node $v$ at layer $l$ is denoted as $h_{v}^{(l)}$. Message passing is performed by the $\text{Propagate}^{(l)}$ function, aggregating information from neighboring nodes of $v$ in layer $l$. The $\text{Aggregate}^{(l)}$ function combines this information with the previous layer’s representation of node $v$ to update $h_{v}^{(l)}$. By incorporating graph structure into learned representations, GNNs can be tailored for tasks like node classification and link prediction.
@@ 3.1. Structural Information Encoding with Text-Graph Grounding
To improve the understanding of graph structural information by large language models, our framework focuses on aligning the encoding of graph structures with the natural language space. This alignment enables language models to effectively comprehend the graph’s structural elements using their language understanding capabilities. To achieve this, we introduce a text-graph grounding paradigm that generates prompts preserving the graph’s structural context for language models. This paradigm acts as a bridge, connecting the semantic understanding of textual information with the inherent structural relationships in the graph.
@@
Text-Structure Alignment. To enhance the alignment of graph structure information with Language Models (LLMs), our focus is on exploring effective encoding methods that can collaborate seamlessly with LLMs. Building upon previous works (Radford et al., 2021; Wen and Fang, 2023), we adopt a contrastive approach by incorporating textual information into the graph structure encoding process. We directly integrate a pre-trained graph encoder into our GraphGPT framework, enabling the seamless utilization of its capabilities. Formally, given a graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$ with raw textual contents $\mathbf{C}={c_{i}\in\mathbb{R}^{l_{i}\times d},1\leq i\leq N}$ for $N$ nodes, we obtain encoded graph representations $\hat{\mathbf{H}}\in\mathbb{R}^{N\times d}$ and encoded text representations $\hat{\mathbf{T}}\in\mathbb{R}^{N\times d}$ as follows:
@@
We utilize the graph encoder, $f_{\mathbf{G}}$, to generate structure-level graph representations from the input graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$. To encode the raw textual contents $\mathbf{C}$ associated with the nodes, we employ a text encoder, such as a transformer or BERT, denoted as $f_{\mathbf{T}}$. This step produces encoded text representations of nodes, which are then normalized row-wise using the norm function. The text-structure alignment across modalities is conducted as follows:
$\Gamma_{1}=(\hat{\mathbf{H}}\hat{\mathbf{T}}^{\top})\cdot\exp(\tau)$ $,\Gamma_{2}=(\hat{\mathbf{H}}\hat{\mathbf{T}}^{\prime\top})\cdot\exp(\tau),\Gamma_{3}=(\hat{\mathbf{T}}^{\top}\hat{\mathbf{T}}^{\prime\top})\cdot\exp(\tau)$
(4) $\mathcal{L}$ $=\sum_{i=1}^{3}\frac{1}{2}\lambda_{i}(\text{CE}(\Gamma_{i},\mathbf{y})+\text{CE}(\Gamma_{i}^{\top},\mathbf{y}))$
where $\hat{\mathbf{T}}^{\prime}=\{\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\hat{\mathbf{T}}_{j},1\leq i\leq N\}$ and $N$ is the number of nodes. In our text-graph grounding, we use the label $\mathbf{y}=(0,1,\cdots,n-1)^{\top}$ for the contrastive alignment objective. We employ a graph transformer (Yun et al., 2019b) as the graph encoder and a vanilla transformer (Vaswani et al., 2017) as the text encoder.
@@ 3.2.1. Self-Supervised Instruction Tuning
In the initial stage of our graph instruction tuning, we introduce self-supervised instruction tuning. This mechanism enhances the language model’s reasoning abilities by incorporating graph domain-specific structural knowledge and effectively understanding contextual information within the graph’s structure. To achieve this, we utilize self-supervised signals derived from unlabeled graph structures as instructions for model tuning. Specifically, we design a structure-aware graph matching task that guides the language model in differentiating between graph tokens using language tokens. This instruction task plays a vital role in accurately associating graph tokens with their corresponding textual descriptions, deepening the model’s comprehension of the graph with the provided guidance.
Instruction Design. The instruction for our graph matching task consists of three components: i) graph information, ii) human question, and iii) GraphGPT response. In this task, we treat each node in the graph as a central node and perform h-hops with random neighbor sampling, resulting in a subgraph structure. The natural language input for the LLM is the human question. In the context of the graph matching task, the instruction includes the indicator token <graph> and a shuffled list of node text information. For example, in a citation graph, the node text information corresponds to paper titles. The objective of the LLM in the graph matching task is to align each graph token with its corresponding node text information. This requires reordering the node text information list based on the sequence of graph tokens, effectively associating each graph token with its relevant textual description. The detailed designs of graph matching are shown in Figure 4.
@@
(5) $p(\mathbf{X}_{\mathcal{O}}|\mathbf{X}_{\mathcal{G}},\mathbf{X}_{\mathcal{I}})=\prod_{i=1}^{L}p_{\theta}(x_{i}|\mathbf{X}_{\mathcal{G}},\mathbf{X}_{\mathcal{I},<i},\mathbf{X}_{\mathcal{O},<i})$
where $\theta$ are the learnable parameters within GraphGPT.
@@ 3.2.2. Task-Specific Instruction Tuning
In the second stage, we introduce task-specific instruction tuning to customize the model’s reasoning behavior for different graph learning tasks, such as node classification or link prediction. By fine-tuning the LLM using task-specific graph instructions, we guide the model to generate responses that align with the constraints and requirements of the specific graph learning task. This enhances the model’s adaptability and performance in handling diverse graph learning tasks.
Figure 3. Workflow of text-structure alignment.
Instruction Design. We utilize a consistent instruction template comprising three parts. To generate graph information for each node, we employ the same neighbor sampling approach as in the first stage. This approach ensures the inclusion of relevant graph information, with each node serving as the central node. For the node classification task, the human question instruction includes the indicator token <graph> and specific text information about the central node. This instruction guides the language model to predict the category of the central node based on both the graph structure data and the accompanying text information. Figure 4 provides instruction examples for different tasks, visually illustrating the presentation of the instruction to the language model.
@@ 3.3. Chain-of-Thought (CoT) Distillation
When faced with diverse graph data, language models may encounter unfamiliar patterns and structures, leading to challenges in generating accurate and coherent responses. This is especially true when the number of node classes varies across different types of graph data, causing distribution shift. To address this challenge and enhance accuracy in the presence of distribution shift, it is crucial to equip our GraphGPT with step-by-step reasoning abilities. Thus, we propose incorporating the Chain-of-Thought (COT) technique (Wei et al., 2022b), which explicitly models the flow of thoughts and reasoning steps. By leveraging COT, our language model improves the coherence and consistency of generated text, enabling it to follow a logical progression of ideas and enhance its understanding and reasoning capabilities for the given graph data.
Incorporating the Chain-of-Thought (COT) technique can be challenging due to the influence of model parameter scale (Shridhar et al., 2023). To overcome this, we draw inspiration from previous research (Shridhar et al., 2023) and adopt a distillation approach. By extracting valuable knowledge from a closed-source, powerful language model like ChatGPT (with over 200 billion parameters), we can generate high-quality COT instructions and enhance our model’s COT reasoning capabilities without increasing the parameter count.
COT Distillation Paradigm. Our approach involves designing tailored Chain-of-Thought (COT) prompts for node-specific tasks. For the node classification task in a citation graph, we provide the abstract, paper title, and a task description as input. Using the GPT-3.5 language model (LLM), we incorporate ”Please think about the categorization in a step-by-step manner.” to enable step-by-step reasoning. By engaging in sequential thought, the LLM generates output that includes predictions for node classes and detailed explanations for each prediction. This ensures transparent and comprehensible reasoning and decision-making. To further enhance performance, we integrate the generated COT instruction data with previously designed instructions for task-specific instruction tuning. With the integrated instructions, we proceed with the proposed instruction tuning paradigm.
@@ 4. Evaluation
We conduct experiments to address key research questions:
@@ 4.1.1. Data Descriptions
We evaluate our GraphGPT using three datasets: OGB-arxiv, PubMed, and Cora. The OGB-arxiv dataset (Hu et al., 2020c) represents a directed graph capturing the citation network among computer science arXiv papers indexed by MAG (Wang et al., 2020). Each paper is manually labeled with a research category selected from 40 subject areas. The PubMed dataset (He et al., 2023) consists of 19,717 scientific publications on diabetes from the PubMed database, categorized into Experimental induced diabetes, Type 1 diabetes, and Type 2 diabetes. Additionally, it includes a citation network with 44,338 links. The Cora dataset (Wen and Fang, 2023) comprises 25,120 research papers connected through citations. We use an expanded version with 70 classes, larger than previous versions (Kipf and Welling, 2017).
@@ 4.1.2. Evaluation Protocols
To facilitate comparison across different datasets, we map node features into a unified vector space by encoding raw text information with a pre-trained BERT (Devlin et al., 2019). In our experiments, we partition the Cora and PubMed datasets into training, validation, and testing sets following a 3:1:1 ratio, as described in previous works (He et al., 2023; Wen and Fang, 2023). For the OGB-arxiv data, we adhere to the public split setting (Hu et al., 2020c) with a training, validation, and testing ratio of 6:2:3. To evaluate our model’s performance, we utilize three commonly used metrics: Accuracy and Macro F1 for node classification, and AUC for link prediction.
@@ 4.1.3. Baseline Methods
In our performance comparison, we consider various state-of-the-art methods for comprehensive evaluation. (i) The first category includes MLP, which employs a Multilayer Perceptron for node representation. (ii) The second category comprises representative graph neural encoders, including GraphSAGE (Hamilton et al., 2017), GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), and RevGNN (Li et al., 2021). (iii) The third category focuses on the self-supervised approach DGI (Velickovic et al., 2019) for graph learning. (iv) The fourth category explores knowledge distillation-enhanced GNNs, with GKD (Yang et al., 2022b) and GLNN (Zhang et al., 2022b) as notable methods. (v). The fifth category showcases recently proposed strong graph transformer networks, with NodeFormer (Wu et al., 2023b) and DIFFormer (Wu et al., 2023a) as competitors. (vi) Lastly, we consider open-sourced LLMs, such as Baichuan-7B, vicuna-7B-v1.1, and vicuna-7B-v1.5 as baselines for understanding text-attributed graph data.